OpenStudy (anonymous):

Evaluate fourth root of 9 multiplied by square root of 9 over the fourth root of 9 to the power of 5.

OpenStudy (anonymous):

i give medals :)

OpenStudy (anonymous):

here are the choices 9 to the power of negative 1 over 2 9 to the power of negative 1 over 4 9 92

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @x.curlsss i give medals :) \(\color{blue}{\text{End of Quote}}\) i prefer chocolates

OpenStudy (anonymous):

lol but can you help me

Nnesha (nnesha):

alright \[\huge\rm \frac{ \sqrt[4]{9} \times \sqrt{9} }{ \sqrt[4]{9^5}} \] like this ?

OpenStudy (anonymous):

what does the 4 at the top mean

Nnesha (nnesha):

4th root

OpenStudy (anonymous):

oh ok

Nnesha (nnesha):

alright so you can convert root to exponent exponent rule \[\huge\rm \sqrt[m]{x^n} = x^ \frac{ n }{ m }\]

Nnesha (nnesha):

m= root number n= exponent of base under the root

OpenStudy (anonymous):

9 4/5 right?

Nnesha (nnesha):

n over m not m over n

Nnesha (nnesha):